Two forces acting at a point whose magnitudes are 8N and 15N if the resultant force is 17N find the angle

Grade:10

2 Answers

Xys
6 years ago
Hello,The answer is pretty simple just use vector addition formula64+225+240*cosβ=289→Cosβ=0Therefore β is 90.
Aravind Kannan Rathinasabapathi
6 years ago
Xys is right,
MORE EXPLAINATION..... is given below
According to law of cosines,
let x be the angle between vectors A and B.
Taking magnitudes,
R2=A2+B2+2ABcos x
R-->Resultant vector Magnitude
172=82+152+2*8*15 cos x
289-289=2*8*15*cos x
cos x=0
x=90o